# A Thermodynamic Constraint for the Electronic Structure of Fe-Ni Alloys at Room And High-Temperature

### v1

@@ Abstract
The measurement of the electronic structure of metal alloys is hampered by the in-applicability of conventional quantum oscillation measurements to high-temperatures and alloy scattering. Recent advancements in the study of the electronic contribution to the entropy suggest a path to ground transport properties in alloy equilibrium thermodynamics. Fe-Ni represents an interesting test-case because it exhibits an intricate electronic structure, order-disorder transformations, and a large solid-solution region where equilibrium data can be obtained. Using cluster modeling, the electronic contribution to the entropy can be inferred from high-temperature thermodynamic data. Electronic transport property measurements can be used to independently evaluate the electronic contribution to the entropy. Reconciling these two approaches at high temperature confirms a new path to study electronic structure for metal alloys.
@@ I Introduction
The inability to reach the high-field condition in pure elements, let alone alloys, renders quantum oscillation measurements unable to support the study of the electronic structure of metallic systems as they approach melting and beyond. Our recent study of Cu-Ni [10] showed the application of an irreversible thermodynamic formalism that links the thermodynamic definition of transport properties to the electronic structure. It suggested an opportunity to use thermodynamic integral quantities, like the entropy and enthalpy, to constraint the electronic structure of metals and alloys. This may become a new method for high-temperature study of metallic systems.
@@ II.1 Rolling and Heat Treatment
20 mm x 8 mm x 4 mm samples were cut using a wire EDM from the rectangular ingots produced from suction casting. Samples were then cold-rolled along the direction of the long axis to a minimum of 30 thickness reduction while cold using an N. Ferrara INC hand roller. Rolled samples were then loaded into quartz ampoules that were purged with Ar 3 times before being vacuum sealed at 10-2 mbar. A special Zr-V-Fe gettering alloy was used to clean any remaining oxygen from the ampoule during annealing (SEAS Getters). Homogenizing temperatures varied between 1050 and 1300°C.
@@ II.5 Interpretations of the Electronic Entropy in Multi-carrier Systems
The formalism used herein links such electronic properties and the electronic state entropy entropy. To calculate the state electronic entropy used to fit the sub-band thermopower and electronic distributions, we evaluate the state electronic entropy at any temperature via equation
@@ III.3 Central Atom Model of Fe-Ni Alloys
The fitting values were produced using a non-linear model fit on the data from Ono et al at 1200 K[40], as detailed in Appendix A. Mathematica was used with the function NonLinearModelFit via the simulated annealing methodology and the parameter space was constrained to physical values of the bonding parameters and cluster stoechiometry as:
Figure (8) shows the enthalpy of mixing from the Central Atom Model with the parameters listed in Appendix A, Table (4). It reproduces key trends with composition of the enthalpy of mixing data of Ono[40] at 1200K. It is found that the cluster on the Fe rich side (Fe3Ni) is unnecessary. Rather the Fe-Ni bond energy is positive to provide a positive enthalpy of mixing regime in that region. The Fe-Ni bond energy is negative, and a driving force is observed for the formation of the FeNi3 compound that is assymetric but otherwise negative.
@@ III.4.3 Order-Disorder Transformation of FeNi3
The temperature variation of the thermoelectric power measured for the FeNi3 samples annealed below their ordering temperature (800K) for various durations prior to measurement is shown in Figure (12). The thermopower converges at the transformation temperature, regardless of the initial degree of ordering in the low-temperature phase. We estimate the thermopower of the ordered phase to be -2 V/K just below the phase-transition temperature thanks to the high-temperature slope to project back to the transformation temperature, and estimate the thermopower of the fully disordered solid solution at 798 K to be -8.8 V/K. We there assume the carrier concentration of Galepov and evaluate the electronic entropy change in a manner described in previous work [6, 7]. Our corresponding estimate of the electronic entropy for the order-disorder transition is -0.318 J/mol K.

### v2

@@ Abstract
The measurement of the electronic structure of metal alloys is hampered by the in-applicability of conventional quantum oscillation measurements owing to electron scattering by thermal or alloy disorder. Recent advancements in the study of the electronic contribution to the entropy suggest a path to ground transport properties in alloy equilibrium thermodynamics. Fe-Ni represents an interesting test-case because it exhibits an intricate electronic structure, order-disorder transformations, and a large solid-solution region where equilibrium data can be obtained. Using cluster modeling, the electronic contribution to the entropy can be inferred from high-temperature thermodynamic data. Electronic transport property measurements can be used to independently evaluate the electronic contribution to the entropy. Reconciling these two approaches at high temperature supports this method to study electronic structure for metal alloys.
@@ I Introduction
The inability to reach the high-field condition in pure elements, let alone alloys, renders quantum oscillation measurements unable to support the study of the electronic structure of metallic systems as they approach melting and beyond. While methods like Angle Resolved Photoemission (ARPES) and x-ray photoemission spectroscopy (XPS) also probe the electronic structure and can be used at high-temperature, these techniques are sensitive to the electronic states on the surface of a material and unlike quantum oscillation measurements, may not be indicative of the bulk electronic states of the system. We therefore confine our comparison only to quantum oscillation measurements when discussing limitations in the prior art.
Our recent study of Cu-Ni [10] showed the application of an irreversible thermodynamic formalism that links the thermodynamic definition of transport properties to the electronic structure. It suggested an opportunity to use thermodynamic integral quantities, like the entropy and enthalpy, to constraint the electronic structure of metals and alloys. This may become a new method for high-temperature study of metallic systems.
@@
Previous work has found that the relative importance of the vibrational entropy is greater for allotropic transitions ([6, 7]) than for alloy mixing ([10]). Therefore, while the electronic contribution is the focus for analysis of solid solution mixing, the vibrational contribution will be examined in the context of the allotropic transition. We do however acknowledge that there is work that may suggest a large role for the vibrational entropy, and that efforts are ongoing to characterize this contribution [19]. Because our arm is to link the electronic contribution to cluster configurational and EMF thermodynamics, rather than compute phase diagrams, it is not within the objective of our current work to address the vibrational contribution in solid solutions.
@@ II.1 Rolling and Heat Treatment
20 mm x 8 mm x 4 mm samples were cut using a wire EDM from the rectangular ingots produced from suction casting. Samples were then cold-rolled along the direction of the long axis to a minimum of 30 thickness reduction while cold using an N. Ferrara INC hand roller. Rolled samples were then loaded into quartz ampoules that were purged with Ar 3 times before being vacuum sealed at 10-2 mbar. A special Zr-V-Fe gettering alloy was used to clean any remaining oxygen from the ampoule during annealing (SEAS Getters). Homogenizing temperatures varied between 1050 and 1300°C. Annealing was conducted to remove the effects of cold rolling on the transport properties and coarsen the grains in order to decrease the grain-boundary effect on subsequent measurement.
@@ II.5 Interpretations of the Electronic Entropy in Multi-carrier Systems
The formalism used herein links such electronic properties and the electronic state entropy. To calculate the state electronic entropy used to fit the sub-band thermopower and electronic distributions, we evaluate the state electronic entropy at any temperature via equation
@@ III.3 Central Atom Model of Fe-Ni Alloys
The fitting values were produced using a non-linear model fit on the data from Ono et al at 1200 K[41], as detailed in Appendix A. Mathematica was used with using the Levenberg-Marquardt method via the parameter space was constrained to physical values of the bonding parameters and cluster stoichiometry as:
Figure (8) shows the enthalpy of mixing from the Central Atom Model with the parameters listed in Appendix A, Table (4). It reproduces key trends with composition of the enthalpy of mixing data of Ono[41] at 1200K. The Fe-Ni bond energy is negative, and a driving force is observed for the formation of the FeNi3 compound that is asymmetric but otherwise negative.
@@ III.4.3 Order-Disorder Transformation of FeNi3
The temperature variation of the thermoelectric power measured for the FeNi3 samples annealed below their ordering temperature (800K) for various durations prior to measurement is shown in Figure (12). The thermopower converges at the transformation temperature, regardless of the initial degree of ordering in the low-temperature phase. We estimate the thermopower of the ordered phase to be 2 V/K just below the phase-transition temperature thanks to the high-temperature slope to project back to the transformation temperature, and estimate the thermopower of the fully disordered solid solution at 798 K to be -8.8 V/K. We there assume the carrier concentration of Galepov and evaluate the electronic entropy change in a manner described in previous work [6, 7]. Our corresponding estimate of the electronic entropy for the order-disorder transition is -0.318 J/mol K.